OpenStudy (anonymous):

Factor the polynomial x^4+10x^2+16

OpenStudy (kirbykirby):

You can let t = x^2, Now you get: P(t) = t^2 + 10t + 16 which looks like a regular quadratic. So, you factor it normally: You get P(t) = (t+8)(t+2)... you get 8 and 2 since 8*2=16 and 8+2 = 10 Now, re-substitute t = x^2: (x^2+8)(x^2+2)